# SAMG: State-Action-Aware Offline-to-Online Reinforcement Learning with Offline Model Guidance

### v1

@@ A.2. Convergence optimality
$\begin{split}\mathbb{E}[\Gamma_{3}]=&\mathbb{E}\left[\gamma p(s_{k},a_{k})\left(Q^{off}(s_{k+1},a_{k+1})\right.\right.\\ &\left.\left.-Q_{\pi}(s_{k+1},a_{k+1})\right)\right]\\ =&\gamma\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)p(s_{k},a_{k})\\ &\left|Q^{off}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|.\end{split}$
@@
$\begin{split}&\big{\|}\mathbb{E}[\Gamma_{3}]\big{\|}_{\infty}\\ =&\big{\|}\gamma\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)p(s_{k},a_{k})\\ &\left|Q^{off}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|\big{\|}_{\infty}\\ =&\gamma p(s_{k},a_{k})\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\bigg{\{}\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}\\ &t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\left|Q^{off}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|\bigg{\}}\\ \leq&\gamma p(s_{k},a_{k})\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\\ &\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\left[\left|Q^{off}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|\right]\\ =&\gamma p(s_{k},a_{k})\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\left[\left|Q^{off}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|\right].\end{split}$
@@
$\begin{split}\big{\|}\mathbb{E}[\Gamma_{3}]\big{\|}_{\infty}&=\gamma\gamma_{\mathcal{A}}p(s_{k},a_{k})\\ &\max\limits_{s^{\prime},a^{\prime}\in\mathcal{D}}\left[\left|Q^{off}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|\right]\\ &=\gamma\gamma_{\mathcal{A}}p(s_{k},a_{k})\big{\|}\Delta_{N}(s,a)\big{\|}_{\infty}.\end{split}$
where $\mathcal{F}$ denotes the function class of offline algorithm, $10<\gamma_{\mathcal{F}}<$ 1 denotes the convergence index of offline algorithm class $\mathcal{F}$ and $N$ denotes the iterative number of offline pre-training.
@@
$\begin{split}\big{\|}\mathbb{E}[\Gamma_{3}]\big{\|}_{\infty}\leq&\max\left\{\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\big{\|}\Delta_{N}(s,a)\big{\|}_{\infty},\right.\\ &\left.2\gamma\xi_{OOD}(s^{\prime},a^{\prime})R_{max}\right\}\\ =&\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\big{\|}\Delta_{N}(s,a)\big{\|}_{\infty}.\end{split}$
@@
$\begin{split}\big{\|}\mathbb{E}[\eta_{k}(s,a)]\big{\|}_{\infty}=&\big{\|}\mathbb{E}[\Gamma_{1}]-\mathbb{E}[\Gamma_{2}]+\mathbb{E}[\Gamma_{3}]\big{\|}_{\infty}\\ \leq&\gamma\left(1-p(s_{k},a_{k})\right)\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}\\ &+\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\big{\|}\Delta_{N}(s,a)\big{\|}_{\infty}.\end{split}$
@@
$\begin{split}\big{\|}\mathbb{E}[\eta_{k}(s,a)]\big{\|}_{\infty}\leq&\gamma\left(1-p(s_{k},a_{k})\right)\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}\\ &+\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\mathcal{O}(\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}).\end{split}$
@@
$\begin{split}&var\left[\eta_{k}(s)|\mathcal{H}_{k}\right]\\ =&var\left\{\left[r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q_{\pi}(s,a)\right]\right.\\ &+\left.\gamma p(s,a)\left(Q^{off}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\right)\right\}.\end{split}$
@@
Appendix B B. Distribution-aware weight
B.1. C-VAE details
B.1.1. Posterior collapse situation
As stated in Section SAMG: methodology, C-VAE may meet with posterior collapse shown in Figure 6. Though the loss seems to be lower in the posterior collapse situation, the output of all state-action samples are 0 and 1 for mean and standard respectively and the encoder totally fail to function.
@@ B.1.1. Posterior collapse situation
B.1.2. VAE implementations
For the C-VAE module, we employ the same VAE structure as Xu [Xu, Zhan, and Zhu, 2022] except that we change the input to (state, action) and the output to next state. Furthermore, we adopt the KL-annealing technique in the hopper environments where we do not introduce the KL loss initially by manually setting it to zero and slowly increasing the KL loss weight with time. KL-annealing could result in more abundant representations of the encoder and is less likely to introduce posterior collapse. We also simplify the decoder of the C-VAE module in hopper and walker2d environments to avoid posterior collapse. Notably, avoid to normalize the states and the actions because the normalized states are highly likely to result in the posterior collapse. In terms of experimental experience, the algorithm performs best when the KL loss converges to around 0.03. The information of next state is supplemented in the training phase to better model the offline distribution and statistical techniques are combined with neural networks to obtain more reasonable probability estimation.
B.1.3. Practical VAE distribution
In practice implementation, the offline data is the input to C-VAE model, and the statistical result of output from the C-VAE model, including mean and standard, are shown in the Figure 7.
@@ B.1.3. Practical VAE distribution
A.2. Embedding network
@@
Appendix C C. Algorithm Implementation
C.1. SAMG implementation on baselines
@@
1:offline Q-network $Q_{\phi}^{off}$, policy $\pi_{\theta}^{off}$ and VAE
2:$\pi_{\theta}\leftarrow\pi_{\theta}^{off}$ , $Q_{\phi}\leftarrow Q_{\phi}^{off}$
3:Initialize the replay buffer $D$ with $N$ samples collected by $Q_{\phi}$
4:for iteration $i=1,2,...$ do
5:     for every environment step do
6:         $a_{t}\sim\pi_{\phi}(s_{t}|s_{t})$
7:         $s_{t+1},d_{t}\sim p_{(}s_{t+1}|s_{t},a_{t})$
8:         insert $(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1},d_{t})$ into D
9:     end for
10:     for every update step do
11:         get $Q_{target}$ and $Q_{target}^{off}$ according to AWAC
12:         get $p_{om}$ with C-VAE according to Section Distribution-aware weight
13:         $Q_{target}\leftarrow(1-p_{om})Q_{target}+p_{om}Q_{target}^{off}$
14:         Update $\phi$ according to Eq. 9 in [Nair et al., 2020] with $Q_{target}$
15:         Update $\theta$ according to Eq. 13 in [Nair et al., 2020]
16:     end for
17:end for
@@
Only some minimal adjustments are needed to implement SAMG on AWAC, and IQL [Kostrikov, Nair, and Levine, 2022] as well. We just need to maintain a much smaller replay buffer filled with online samples and insert and sample from this "online replay buffer". Before conducting normal gradient update step, we need to calculate the mixed $Q_{target}$ according to 4. As for IQL, we freeze and query the offline pre-trained value function instead because IQL separately trains a value function to serve as the target information. Other implementations are similar to AWAC and are omitted.
@@
As for CQL and Cal_QL, these two algorithms share a similar implementation procedure and align with AWAC when calculating the $Q_{target}$. However, CQL adds one extra penalty term to minimize the expected Q-value based on a distribution $\mu(a|s)$, formulated as $\mathbb{E}_{s\sim\mathcal{D},a\sim\mu(a|s)}\left[Q(s,a)\right]$. This term is separated from the standard Bellman equation and serves an important role in making sure the learned Q-function is lower-bounded. However, this term is unrestricted in our paradigm and may cause algorithm divergence. So we add an offline version of the term still weighted by the distribution-aware weight. This slightly avoids our setting but is reasonable that this setting shares the consistent updating direction with the Bellman equation error term.
@@
We implement all the algorithms based on the benchmark CORL [Tarasov et al., 2024], whose open source code is available at https://github.com/tinkoff-ai/CORL and the license is Apache License 2.0 with detail in the GitHub link.
C.2. Implementation Details
@@
In practice, we strictly adopt the CORL setting to train the offline model and the vanilla fine-tuning training, including the training process and hyperparameters. In the fine-tuning process, training for all algorithms other than IQL is conducted for 500,000 steps, while IQL is conducted for 1,000,000 steps. As for SAMG training, for mujoco environments (halfcheetah, hopper, walker2d), SAMG algorithms share the same set of hyperparameters with the fine-tuning process to illustrate fairness. In the antmaze environment, we slightly reduce the weight of the Q-value maximization term to highlight the impact of SAMG for algorithms CQL and Cal_QL, from 5 to 2-3. For the threshold $p_{om}$, we have shown in Section Sensitivity analysis of threshold $p_{om}$ that it is not sensitive in most environments, we adopt the value of 0.5 in most environments which seems large but only a small portion satisfies the condition. For the antmaze environment, we take 0.4 for Cal_QL and 0.5 for CQL. We maintain an online replay buffer with a small capacity (only 50,000) and initialize it with 1000 samples utilizing the offline model (1000 is the normal length of an episode in most environments). The details of C-VAE are jointly stated in Appendix B.
C.3. Datasets

### v2

@@ A.2 Convergence optimality
$\begin{split}\mathbb{E}[\Gamma_{3}]=&\mathbb{E}\left[\gamma p(s_{k},a_{k})\left(Q^{off}(s_{k+1},a_{k+1})-Q^{\pi}(s_{k+1},a_{k+1})\right)\right]\\ =&\gamma\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)p(s_{k},a_{k})\left|Q^{off}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\right|.\end{split}$
@@
$\begin{split}&\big{\|}\mathbb{E}[\Gamma_{3}]\big{\|}_{\infty}\\ =&\Big{\|}\gamma\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)p(s_{k},a_{k})\bigl{|}Q^{off}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\bigr{|}\Big{\|}_{\infty}\\ =&\gamma p(s_{k},a_{k})\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\bigg{\{}\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\bigl{|}Q^{off}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\bigr{|}\bigg{\}}\\ \leq&\gamma p(s_{k},a_{k})\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\Bigl{[}\bigl{|}Q^{off}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\bigr{|}\Bigr{]}\\ =&\gamma p(s_{k},a_{k})\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\Bigl{[}\bigl{|}Q^{off}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\bigr{|}\Bigr{]}.\end{split}$
@@
$\begin{split}\big{\|}\mathbb{E}[\Gamma_{3}]\big{\|}_{\infty}&=\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\max\limits_{s^{\prime},a^{\prime}\in\mathcal{D}}\Bigl{[}\bigl{|}Q^{off}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\bigr{|}\Bigr{]}\\ &=\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\big{\|}\Delta_{N}(s,a)\big{\|}_{\infty}.\end{split}$
where $\mathcal{F}$ denotes the function class of offline algorithm, $10<\gamma_{\mathcal{F}}<$ 1 denotes the convergence coefficient of offline algorithm class $\mathcal{F}$ and $N$ denotes the iterative number of offline pre-training.
@@
$\begin{split}\big{\|}\mathbb{E}[\Gamma_{3}]\big{\|}_{\infty}\leq&\max\Bigl{\{}\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\big{\|}\Delta_{N}(s,a)\big{\|}_{\infty},2\gamma\xi_{OOD}(s^{\prime},a^{\prime})R_{max}\Bigr{\}}\\ =&\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\big{\|}\Delta_{N}(s,a)\big{\|}_{\infty}.\end{split}$
@@
$\begin{split}\big{\|}\mathbb{E}[\eta_{k}(s,a)]\big{\|}_{\infty}=&\big{\|}\mathbb{E}[\Gamma_{1}]-\mathbb{E}[\Gamma_{2}]+\mathbb{E}[\Gamma_{3}]\big{\|}_{\infty}\\ \leq&\gamma\left(1-p(s_{k},a_{k})\right)\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}+\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\big{\|}\Delta_{N}(s,a)\big{\|}_{\infty}.\end{split}$
@@
$\begin{split}\big{\|}\mathbb{E}[\eta_{k}(s,a)]\big{\|}_{\infty}\leq&\gamma\left(1-p(s_{k},a_{k})\right)\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}+\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\mathcal{O}(\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}.\end{split}$ (18)
@@
$\begin{split}&var\bigl{[}\eta_{k}(s)|\mathcal{H}_{k}\bigr{]}\\ =&var\Bigl{\{}\bigl{[}r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q^{\pi}(s,a)\bigr{]}+\gamma p(s,a)\bigl{[}Q^{off}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\bigr{]}\Bigr{\}}.\end{split}$
@@
A.3 Convergence Speed
In this section, we give a more detailed analysis of the convergence speed of SAMG. For vanilla RL algorithms, the contraction coefficient $\gamma$ represents the convergence speed because it controls the contraction speed of Q-iteration. For SAMG, we give a rough derivation in Appendix A.1 that SAMG possesses a smaller contraction coefficient. But how small could that be? We actually have already derived the specific form of contraction factor in Appendix A.2, as specified in Eq. (18). However, Eq. (18) just covers the in-distribution situation of the contraction coefficient. As for the OOD situation, the contraction coefficient share the same coefficient as the normal Bellman equation. To sum up, we write the whole the contraction coefficient as below:
$\gamma_{o}\leq\left\{\begin{aligned} &\begin{aligned} &\left(1-p(s_{k},a_{k})\right)\gamma+\gamma\gamma_{\mathcal{F}}p(s_{k},a_{k})\frac{\big{\|}\Delta_{off}(s,a)\big{\|}_{\infty}}{\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}},\end{aligned}&p(s_{k},a_{k})\geq p_{m}\\ &\gamma,&p(s_{k},a_{k})<p_{m}\end{aligned}\right.$ (19)
where $10<\gamma_{\mathcal{F}}<$ 1 denotes the convergence coefficient of offline algorithm class $\mathcal{F}$ and $\big{\|}\Delta_{off}(s,a)\big{\|}_{\infty}$ denotes the offline suboptimality bound $\left[V^{*}(s)-V^{\pi_{off}}(s)\right]$ and $\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}$ denotes the suboptimality bound of the k-th iteration of online fine-tuning.
The upper equation holds for in-distribution samples, which are well mastered by the offline model. Therefore, the offline suboptimality bound is substantially tighter compared to the online bound. This illustrates that the offline model guidance significantly accelerates the online fine-tuning process by providing accurate estimations for in-distribution samples.
Appendix B State-action-conditional coefficient
B.1 C-VAE details
B.1.1 Posterior collapse situation
As stated in Section 3, C-VAE may meet with posterior collapse, as shown in Fig. 6. Fig. 6 illustrates the KL loss values with posterior collapse (with “s-hopper-m" and “s-hopper-me" legend) and without posterior collapse (with “sa-hopper-m" and “sa-hopper-me" legend), representing the distribution error between the output distribution of encoder and standard normal distribution. It can be observed that situations with posterior collapse possess much lower loss term (approximately by four orders of magnitude). Though the loss is lower for the posterior collapse situation, the output of all state-action samples are 0 and 1 for mean and standard respectively. Therefore the encoder totally fails to function.
@@ B.1.1 Posterior collapse situation
B.1.2 VAE implementations
For the C-VAE module, we employ the same VAE structure as Xu [Xu et al., 2022] except that we change the input to (state, action) and the output to next state. Furthermore, we adopt the KL-annealing technique in the hopper environments where we do not introduce the KL loss initially by manually setting it to zero and slowly increasing the KL loss weight with time. KL-annealing could result in more abundant representations of the encoder and is less likely to introduce posterior collapse. We also simplify the decoder of the C-VAE module in hopper and Walker2d environments to avoid posterior collapse. Notably, avoid normalizing the states and the actions because the normalized states are highly likely to result in the posterior collapse. In terms of experimental experience, the algorithm performs best when the KL loss converges to around 0.03. The information of the next state is supplemented in the training phase to better model the offline distribution and statistical techniques are combined with neural networks to obtain more reasonable probability estimation.
B.1.3 Practical VAE distribution
In practice implementation, the offline data is the input to C-VAE model, and the statistical result of output from the C-VAE model, including mean and standard, are shown in the Fig. 7.
@@ B.1.3 Practical VAE distribution
B.1.4 Adaptive VAE Coefficient
By adding a weight variable to the replay buffer, we store the offline critic value and $p^{off}$ for each sample. At fixed intervals, we first collect data from the current period (all data from the previous to the current interval) and then filter out OOD samples, whose $p^{off}$ is equal to 0. The challenge lies in identifying the mastered OOD samples. Since the model is lack of awareness of OOD samples, estimated Q-values tend to introduce significant errors and, consequently, large loss terms during training. This can destroy the model and reduce algorithm performance. Therefore, the magnitude of the error between the estimated critic during online fine-tuning and the true Q-values can be used as a measure of how well the OOD samples have been mastered. However, in practice, the ground truth of the Q-values is unavailable, so the exact error can not be obtained. To address this, we adopted several potential approaches to estimate the error, which will be detailed later. After obtaining error estimates, we select samples with minimal errors as mastered OOD samples. These samples are then used to continue training the VAE model to help it learn these samples.
In the actual implementation, the error estimation methods replace the true Q-values with sampling (as referenced in [Kumar et al., 2020]) and use the practical Bellman operator, where the target Q-value serves as an estimate. We found that the results of these two methods are similar, with over 70% overlap in the filtered OOD samples, and both lead to comparable improvements in algorithm performance (with the sampling-based method performing slightly better). Considering the trade-off between the computation overhead and algorithm performance, we choose to use the Bellman operator. Additionally, we set the update interval to 10,000 steps. Since this interval exceeds the target Q network update interval (typically set to 1,000), we consistently refer to the target-Q network at the beginning of the period to ensure fairness. Given that the target network is already saved at this point, no additional computation overhead is introduced. Minimal error is defined as the smallest 10% of errors among the filtered samples.
B.2 Embedding network
@@
Appendix C Algorithm Implementation
C.1 SAMG implementation on baselines
@@
Algorithm 1 Offline-to-Online Reinforcement Learning via State-Action-Conditional Offline Model Guidance (Implemented on AWAC)
0:  offline Q-network $Q_{\phi}^{off}$, policy $\pi_{\theta}^{off}$ and VAE
1:  $\pi_{\theta}\leftarrow\pi_{\theta}^{off}$ , $Q_{\phi}\leftarrow Q_{\phi}^{off}$
2:  Initialize the replay buffer $D$ with $N$ samples collected by $Q_{\phi}$
3:  for iteration $i=1,2,...$ do
4:     for every environment step do
5:        $a_{t}\sim\pi_{\phi}(s_{t}|s_{t})$
6:        $s_{t+1},d_{t}\sim p_{(}s_{t+1}|s_{t},a_{t})$
7:        insert $(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1},d_{t})$ into D
8:     end for
9:     for every update step do
10:        get $Q_{target}$ and $Q_{target}^{off}$ according to AWAC
11:        get $p_{om}$ with C-VAE according to Section 3.2
12:        $Q_{target}\leftarrow(1-p_{om})Q_{target}+p_{om}Q_{target}^{off}$
13:        Update $\phi$ according to Eq. 9 in [Nair et al., 2020] with $Q_{target}$
14:        Update $\theta$ according to Eq. 13 in [Nair et al., 2020]
15:        if This step is the interval to update coefficient then
16:           Filter mastered OOD samples
17:           Train the VAE model with collected samples
18:           Update the VAE model and the offline critic
19:        end if
20:     end for
21:  end for
@@
Only some minimal adjustments are needed to implement SAMG on AWAC, and IQL [Kostrikov et al., 2022] as well. We just need to maintain a much smaller replay buffer filled with online samples and insert and sample from this “online replay buffer". Before conducting normal gradient update step, we need to calculate the mixed $Q_{target}$ according to Eq. (5). As for IQL, we freeze and query the offline pre-trained value function instead because IQL separately trains a value function to serve as the target information. Other implementations are similar to AWAC and are omitted.
@@
As for CQL and Cal_QL, these two algorithms share a similar implementation procedure and align with AWAC when calculating the $Q_{target}$. However, CQL adds one extra penalty term to minimize the expected Q-value based on a distribution $\mu(a|s)$, formulated as $\mathbb{E}_{s\sim\mathcal{D},a\sim\mu(a|s)}\left[Q(s,a)\right]$. This term is separated from the standard Bellman equation and serves an important role in making sure the learned Q-function is lower-bounded. However, this term is unrestricted in our paradigm and may cause algorithm divergence. So we add an offline version of the term still weighted by the state-action-conditional coefficient. This slightly avoids our setting but is reasonable that this setting shares the consistent updating direction with the Bellman equation error term.
@@
We implement all the algorithms based on the benchmark CORL [Tarasov et al., 2024], whose open source code is available at https://github.com/tinkoff-ai/CORL and the license is Apache License 2.0 with detail in the GitHub link. Our code is attached in the supplementary material.
C.2 Implementation Details
@@
In practice, we strictly adopt the CORL setting to train the offline model and the vanilla fine-tuning training, including the training process and hyperparameters. As for SAMG training, for mujoco environments (halfcheetah, hopper, walker2d), SAMG algorithms share the same set of hyperparameters with the fine-tuning process to illustrate fairness. In the antmaze environment, we slightly reduce the weight of the Q-value maximization term to highlight the impact of SAMG for algorithms CQL and Cal_QL, from 5 to 2. For the threshold $p^{vae}_{m}$, we adopt the value of 0.6 in most environments (including HalfCheetah, Hopper, Walker2d and Adroit) which seems large but only a small portion satisfies the condition. For the antmaze environment, we take 0.7 for CQL and 0.6 for the others. We found that in our setting, reducing the size of the replay buffer allows for more efficient utilization of samples, thereby improving the algorithm’s performance. Specifically, we set the buffer size to be 50,000 for Anrmaze environment and 20,000 for the other environments. We initialize the replay buffer with 2000 samples utilizing the offline model (2000 is the normal length of an episode in most environments). The details of C-VAE have been stated in Appendix B.
C.3 Datasets